OpenStudy (anonymous):

HELP ILL GIVE MEDALS f-1(3) when f(x) = (2x-7)/ 3

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

answer this question by solving \[\frac{2x-7}{3}=3\] for (x\)

OpenStudy (misty1212):

you ok with doing that or you need help?

OpenStudy (anonymous):

alright let me think..

OpenStudy (misty1212):

ok i'll wait the first step is to multiply by 3

OpenStudy (anonymous):

2x-7/3 * 3/1 = 2x-7=3?

OpenStudy (misty1212):

\[\frac{2x-7}{3}=3\\ 2x-7=9\]

OpenStudy (anonymous):

Oh okay because you have to multiply both sides?

OpenStudy (anonymous):

2x-7=9 2x=16 x=8?

OpenStudy (misty1212):

looks good to me

OpenStudy (anonymous):

Thanks so much!

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]